OpenStudy (anonymous):

PLEASE CHECK IF MY ANSWERS CORRECT The residents of a city voted on whether to raise property taxes. The ratio of yes votes to no votes was 5 to 8. If there were 2660 yes votes, what was the total number of votes?

OpenStudy (anonymous):

am sorry don't know

OpenStudy (anonymous):

\[\frac{ 5 }{ 8 } = \frac{ 2660 }{ x }\] \[5x = 2660 \times 8 \] \[x = \frac{ 2660 \times 8 }{ 5}\] \[x = 4256\]

OpenStudy (anonymous):

one min.

OpenStudy (anonymous):

nevermind i got the answer it was 6916

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

ratio of yes votes to no votes = 5 : 8 total number of parts = 5 + 8 = 13 Number of yes votes= 2660 (5 parts out of 13) Let x be the number of total votes. 5/13 of x = 2660 (5/13)x = 2660 5x = 2660 * 13 x = (2660 * 13)/5 x = 6916 Hence the number of total votes is 6916. sorry got a bit busy :)

OpenStudy (anonymous):

Lol you already got the answer :D